# LeGrad: An Explainability Method for Vision Transformers via Feature Formation Sensitivity

### v1

@@ 4.2 Comparison to State-of-the-Art
Table 3: SOTA Perturbation Performance: Comparison of interpretability methods on the ImageNet-val using a different model size.
ImageNet OpenImagesV7
Negative Positive
rollout 47.81 47.81 25.74 25.74 0.07
Raw attention 44.42 44.42 25.85 25.85 0.09
GradCAM 41.25 44.42 35.10 33.50 6.97
AttentionCAM 45.62 45.71 45.01 44.92 0.19
CheferCAM 47.12 49.13 22.35 21.15 1.94
LeGrad 50.08 51.67 18.48 17.55 25.40
Table 4: SOTA comparison on SigLIP-B/16: Comparison of interpretability methods on perturbation-based tasks on ImageNet-val and open-vocabulary localization on OpenImagesV7.
4.3 Ablation
4.3.1 Ablation Studies on Layer Contributions
@@
Ablation studies are conducted to understand the impact of the number of layers considered in the computation of LeGrad’s explainability maps. We investigate this aspect across two distinct benchmarks: a perturbation-based evaluation using the ImageNet validation set and an open-vocabulary segmentation task on the OpenImagesV7 dataset. The experiments are performed on various model sizes from the openCLIP library, including ViT-B/16, ViT-L/14, and ViT-H/14.
@@
Perturbation-Based Evaluation: In the perturbation-based evaluation, see Figure 4(a), we employ a negative perturbation test as described in Section 4.1, using the ground-truth class for reference. The results indicate that the ViT-B/16 model’s performance is optimal when fewer layers are included in the explainability map computation. Conversely, larger models such as ViT-L/14 and ViT-H/14 show improved performance with the inclusion of more layers. This suggests that in larger models, the aggregation of information into the [CLS] token is distributed across a greater number of layers, necessitating a more comprehensive layer-wise analysis for accurate explainability.
@@
Open-Vocabulary Segmentation: For the open-vocabulary segmentation task, Figure 4(b), all models demonstrate enhanced performance with the inclusion of additional layers in the explainability map computation. The optimal number of layers varies with the model size, with larger models requiring a greater number of layers. This finding aligns with existing literature [18], which suggests that the information aggregation process in ViTs is more distributed in larger architectures. However, it is also observed that beyond a certain number of layers, the performance gains plateau, indicating that the inclusion of additional layers does not further enhance the explainability map’s quality.
These ablation studies underscore the importance of layer integration in the construction of explainability maps for Vision Transformers. They further show that while the inclusion of more layers can be beneficial, especially for larger models, there exists an optimal range beyond which no additional interpretability is gained. Indeed, we note in Figure 4 that including the last $40\%$ of the ViT layers in LeGrad’s explainability computation is a good trade-off. This can be helpful for the practical application of LeGrad, as it allows for a more targeted and computationally efficient approach to generating explainability maps without compromising on their quality.
(a) AUC for Negative perturbation on ImageNet-val for different layers used for LeGrad
(b) point-mIoU on OpenImagesV7 for different layers used for LeGrad.
@@
Figure 4: Ablation on the number of layers used in LeGrad for different architecture sizes.
4.3.2 Ablations of ReLU and Layer Aggregation
L/14 H/14
[…]
47.81 20.80 57.57 21.73
49.32 19.95 59.55 19.50
52.01 16.80 60.28 18.26
54.48 15.23 61.72 18.26
Table 5: Ablation study: "ReLU" corresponds to whether or not negative gradients are set to 0. "All layers" corresponds to whether or not the intermediate tokens are used to compute the gradient for every layer or if only the features from the last layer are used. Numbers are the AUC score for the perturbation base benchmark using the target class to compute the explainability map.
4.4 Qualitative analysis
4.4.1 Ablation Studies on Layer Contributions
To elucidate the interpretability of Vision Transformers (ViTs) through our proposed LeGrad method, we extend our examination beyond quantitative ablation studies to a qualitative analysis. This analysis scrutinizes the localization maps generated by LeGrad for the last ten layers of various ViT models, as well as the final map that merge these individual layer contributions. The experiments are performed on various model sizes from the openCLIP library, including ViT-B/16, ViT-L/14, and ViT-H/14.
The qualitative results, illustrated in Figure 5, reveal that the localization of the prompt is not confined to a single layer but is distributed across multiple layers. This observation is consistent with the findings from the ablation studies in Section 4.3, which underscore the necessity of incorporating multiple layers into the explainability framework to capture the full scope of the model’s decision-making process.
A notable distinction is visible in the behavior of smaller models, such as ViT-B/16, with that of larger counterparts like ViT-L/14 and ViT-H/14. In the case of ViT-B/16, the explainability signal predominantly emanates from the final layers, particularly the last two. Conversely, in larger models, earlier layers also substantially influence the merged explainability map. For instance, in Figure 5, layers 8 and 9 of ViT-H/14 appear to distinctly focus on the heads of the cats, while other layers’ contributions are more diffused. This observation not only corroborates the utility of incorporating multiple layers into the explainability analysis but also suggests a more distributed information aggregation process into the [CLS] token in larger models, as posited in the literature [18].
The qualitative analysis thus reinforces the hypothesis that the mechanism by which ViTs integrate information across layers is size-dependent, with larger models engaging a broader range of layers in the synthesis of the global feature representation. This is important for the practical application of LeGrad, as it informs the selection of layers to be included in the explainability map computation, ensuring both computational efficiency and the retention of fidelity.
@@ 4.4.1 Ablation Studies on Layer Contributions
Figure 5: Qualitative analysis of the impact of each layer for different model sizes using "a photo of a cat" as prompt. In smaller models the explainability signal predominantly emanates from the final layers while in larger models lower layers also contribute to the explainability map.
4.4.2 Qualitative Comparison to SOTA
Here, we present a qualitative analysis of the explainability maps generated by LeGrad in comparison to other state-of-the-art (SOTA) methods. The visual results are depicted in Figure 6, which includes a diverse set of explainability approaches such as gradient-based methods (e.g., CheferCAM, GradCAM), attention-based methods (e.g., Raw Attention weights visualization, Rollout), and methods that integrate intermediate visual representations with text prompts (e.g., TextSpan).
Our observations indicate that raw attention visualizations tend to highlight a few specific pixels with high intensity, often associated with the background rather than the object of interest. This pattern, consistent with findings in the literature [9, 15], suggests that certain tokens disproportionately capture attention weights. Consequently, methods that rely on raw attention weights to construct explainability maps, such as CheferCAM, exhibit similar artifacts. For instance, in the localization of "Wheel" (Figure 6, row 2), the wheel is marginally accentuated amidst a predominance of noisy, irrelevant locations.
In contrast, LeGrad is resilient to such distractions by leveraging gradients with respect to attention maps. The presence of uniform noisy activations across different prompts results in minimal gradients for these regions, effectively filtering them out from the final heatmaps. This characteristic enables LeGrad to produce more focused and relevant visual explanations.
The presence of noisy activations is also observed in methods like TextSpan and GradCAM, which utilize intermediate feature representations. These artifacts tend to surface in the same spatial regions as those highlighted by raw attention weights, suggesting a potential link to tokens with inflated attention. While the underlying cause of this behavior demands further investigation, it is beyond the scope of the current work. Eventually, Figure 6 showcases LeGrad’s capability to accurately localize a wide range of objects and concepts, from commonplace items like "baskets" to abstract notions such as "numbers" and activities like "swimming." The method’s robustness and versatility underscore its effectiveness in providing interpretable insights into the decision-making processes of vision transformers.
@@ 5 Conclusion
In this work, we proposed LeGrad as an explainability method that highlights the decision-making process of Vision Transformers (ViTs) across all layers various feature aggregation strategies. We validated the method’s effectiveness in generating interpretable visual explanations that align with the model’s reasoning. Our approach offers a granular view of feature formation sensitivity, providing insights into the contribution of individual layers and attention mechanisms. The evaluation across object segmentation, perturbation-based metrics, and open-vocabulary scenarios underscores LeGrad’s versatility and fidelity in different contexts. By facilitating a deeper understanding of ViTs LeGrad paves the way for more transparent and interpretable foundation models.
@@
LeGrad: An Explainability Method for Vision Transformers via Feature Formation Sensitivity
@@ 7 Overview
In this supplementary material, we first provide in section 8 a link to a demo of our novel method, LeGrad, showcasing its practical application. Detailed information regarding the implementation, including pre-trained weights and baseline methodologies for comparative evaluation, is presented in section 9. An expanded evaluation of LeGrad’s performance, specifically on Vision Transformers (ViT) trained with the ImageNet dataset, is documented in section 10. Section 10.2 offers a visual representation of the perturbation benchmarks utilized to assess the efficacy of various explainability approaches. Additional qualitative examples are provided in section 10.3 to further illustrate the capabilities of our method. Lastly, section 11 includes a disclaimer addressing the use of personal and human subjects data within our research.
8 Demo
An image is worth a $1000$ words, we, therefore, provide a demo at:
https://huggingface.co/spaces/WalidBouss/LeGrad-Demo.
@@ 9.1 Pretrained weights
The experiments conducted in our study leverage a suite of models with varying capacities, including ViT-B/16, ViT-L/14, ViT-H/14, and ViT-bigG/14. These models are initialized with pretrained weights from the OpenCLIP library respectively identified by: "laion2b_s34b_b88k", "laion2b_s32b_b82k", "laion2b_s32b_b79k", and "laion2b_s39b_b160k". For the SigLIP method, we utilize the ViT-B/16 model equipped with the "webl" weights.
@@ 9.2 Detailed Description of Baselines
In this section, we provide a concise overview of the baseline methods against which our proposed approach is compared.
@@
with $d$ representing the token dimension, * denoting element-wise multiplication, and the superscript $+$ indicating the ReLU operation. We empirically determined that applying GradCAM to layer 8 of ViT-B/16 yields optimal results.
@@ 9.3 Adaptation of Baseline Methods to Attentional Pooler
CheferCAM: Following the original paper [13] that introduces CheferCAM, we considered the attentional pooler as an instance of a "decoder transformer" and applied the relevancy update rules described in equation (10) of that paper [13], (following the example of DETR). Since the attentional pooler has no skip connection we adapted the relevancy update rule not to consider the skip connection.
@@
9.4 Mitigation of sensitivity to irrelevant regions:
$E^{s}_{E^{empty}>th}=0,$ (12)
where $th=0.8$. This method effectively addresses the issue without resorting to external data regarding the image content.
10 Additional results
10.1 Image Classification
In this section, we extend our evaluation of the proposed LeGrad method to Vision Transformers (ViTs) that have been trained on the ImageNet dataset for the task of image classification. The results of this evaluation are presented in Table 7, also providing a comparison with other state-of-the-art interpretability methods.
Another observation is that even elementary interpretability approaches, such as examining the raw attention maps from the final attention layer of the ViT, demonstrate competitive results. In fact, these basic methods surpass more complex ones like GradCAM (achieving an AUC of 53.1 versus 43.0 for negative perturbations).
@@ 10.1 Image Classification
Table 7: SOTA comparison on ViT-B/16: Comparison of interpretability methods on perturbation-based tasks on ImageNet-val for a ViT trained on ImageNet.

### v2

@@ 4.2 Open-Vocabulary localization
rollout 47.81 47.81 25.74 25.74 0.07
Raw attention 44.42 44.42 25.85 25.85 0.09
GradCAM 41.25 44.42 35.10 33.50 6.97
AttentionCAM 45.62 45.71 45.01 44.92 0.19
CheferCAM 47.12 49.13 22.35 21.15 1.94
LeGrad 50.08 51.67 18.48 17.55 25.40
Table 4: SOTA comparison on SigLIP-B/16: Comparison of explainability methods on perturbation-based tasks on ImageNet-val and open-vocabulary localization on OpenImagesV7.
Method Speech Seg. Sound Seg.
[…]
DAVENet 32.2 26.3 16.8 17.0
CAVMAE 27.2 19.9 26.0 20.5
ImageBind 20.2 19.7 18.3 18.1
ImageBind + LeGrad 23.3 21.8 48.0 38.9
DenseAV* 48.7 36.8 32.7 24.2
Table 5: Speech & Sound prompted semantic segmentation: Comparison of the sound localization methods on ADE20K Speech & Sound Prompted dataset [19].
4.3 Audio Localization
[…]
Table 5 compares the sound localization performance of different state-of-the-art methods. We observe that when applied to the ImageBind model [16], LeGrad leads to a significant increase in sound segmentation performance, hence establishing a new SOTA on that benchmark by outperforming DAVENet [20], CAVMAE [18] and the previous SOTA DenseAV [19]. The less pronounced improvement of LeGrad over ImageBind on speech segmentation is due to the fact that ImageBind was not trained with speech data (only sound). Therefore, since ImageBind performs poorly on speech, LeGrad cannot improve the localization further.
4.4 Perturbation-Based Evaluation
[…]
As detailed in Table 6, LeGrad’s performance is here comparable to TextSpan for positive perturbations and slightly superior for negative perturbations. For all other methods, LeGrad outperformes both attention-based (e.g., ”rollout” and ”raw attention”) and gradient-based methods (e.g., GradCAM, AttentionCAM, and CheferCAM) across various model sizes and for both predicted and ground truth classes, emphasizing its ability to identify and preserve critical image regions for accurate classification.
4.5 Speed Comparison
Table 3 compares the inference speed for different methods averaged over $1,000$ images. Generally, gradient-based methods are faster than methods like LRP and TextSpan. More specifically, despite using several layers for its prediction, LeGrads speed only drops slightly compared to GradCAM and AttentionCAM, which both use a single layer and are significantly faster than CheferCAM. The speed difference stems from summing the contribution of each layer rather than using matrix multiplication as in CheferCAM.
4.6 Performance on SigLIP
We also evaluate the adaptability regarding the performance on SigLIP-B/16, a Vision-Language model employing an attentional pooler as shown in Table 5. The results underscore the methods performance across both negative and positive perturbation-based benchmarks. Notably, in the open-vocabulary benchmark on OpenImagesV7, LeGrad achieved a p-mIoU of 25.4, significantly surpassing GradCAM’s 7.0 p-mIoU, the next best method. These findings affirm the versatility of LeGrad, demonstrating its robust applicability to various pooling mechanisms within Vision Transformers. Further details on the methodological adaptations of LeGrad and other evaluated methods for compatibility with SigLIP are provided in the annex.
Negative Positive
Method Pred. $\uparrow$ Targ. $\uparrow$ Pred. $\downarrow$ Targ. $\downarrow$
@@ 4.6 Performance on SigLIP
Method Pred. $\uparrow$ Targ. $\uparrow$ Pred. $\downarrow$ Targ. $\downarrow$
@@
Table 6: SOTA Perturbation Performance: Comparison of explainability methods on the ImageNet-val using a different model size.
4.7 Ablation Studies
@@
Layer Accumulation.To further understand the impact of the number of layers considered in the computation of LeGrad’s explainability maps, we investigate this aspect across two distinct benchmarks: a perturbation-based evaluation using the ImageNet validation set and an open-vocabulary segmentation task on the OpenImagesV7 dataset. The experiments are performed on various model sizes from the OpenCLIP library, including ViT-B/16, ViT-L/14, and ViT-H/14.
@@
$\diamondsuit$ Perturbation-Based Evaluation: In the perturbation-based evaluation (Figure 4 left) we employ a negative perturbation test as described in Section 4.4, using the ground-truth class for reference. The results indicate that the ViT-B/16 model’s performance is optimal when fewer layers are included in the explainability map computation. Conversely, larger models such as ViT-L/14 and ViT-H/14 show improved performance with the inclusion of more layers. This suggests that in larger models, the aggregation of information into the [CLS] token is distributed across a greater number of layers, necessitating a more comprehensive layer-wise analysis for accurate explainability.
@@
$\diamondsuit$ Open-Vocabulary Segmentation: For the open-vocabulary segmentation task (Figure 4 right) all models demonstrate enhanced performance with the inclusion of additional layers in the explainability map computation. The optimal number of layers varies with the size of the model, with larger models requiring a larger number of layers. This finding is consistent with existing literature [14], suggesting that the information aggregation process in ViTs is more distributed in larger architectures. However, it is also observed that beyond a certain number of layers, the performance plateaus, indicating that the inclusion of additional layers does not further enhance the explainability map’s quality. To further analyze those effects, we provide a qualitative visualization of the per-layer heatmaps produced by LeGrad in Figure 5. It shows that the localization of the prompt is not confined to a single layer but is distributed across multiple layers. This observation is consistent with the findings from the above ablations, which underscore the necessity of incorporating multiple layers into the explainability framework to capture the full scope of the model’s decision-making process. This observation not only corroborates the utility of incorporating multiple layers into the explainability analysis but also suggests a more distributed information aggregation process into the [CLS] token in larger models, as posited in the literature [14].
@@
Figure 4: Ablation on the number of layers used in LeGrad for different architecture sizes. (Up): AUC for Negative perturbation on ImageNet-val for different layers used for LeGrad. (Down): point-mIoU on OpenImagesV7 for different layers used for LeGrad.
@@
Figure 5: Qualitative analysis of the impact of each layer for different model sizes using ”a photo of a cat” as prompt. In smaller models, the explainability signal predominantly emanates from the final layers, while in larger models, lower layers also contribute.
@@
Gradient Distribution Analysis. To better understand the inner workings of the proposed LeGrad method, we further analyse the gradient intensity distribution across the layers of ViT-L/14 models from OpenCLIP based on Laion400M, OpenAI CLIP, and MetaCLIP. This analysis was performed using the PascalVOC dataset. The explainability maps for each layer were computed in accordance with Equation 8. Subsequently, the associated segmentation mask was utilized to compute the average gradient within the mask for each layer. This process yielded a single value per layer. To account for the impact of the mask size, we apply min-max normalization across all layers. This normalization process ensured that the layer contributing the most to the final explainability map was assigned a value of one, while the least contributing layer was assigned a value of zero. Figure 7 depicts the gradient distribution over the layers for different sets of pretrained ViT-L/14 models. First, we observe that for most models, the layers contributing the most are typically located towards the end of the ViT. Second, despite all sharing the exact same model architecture and being trained with the same loss, we observe significant difference in layer importance. For Laion400M and OpenAI the most important layer is the last one whereas for MetaCLIP it is rather the penultimate. Interestingly, for the Laion400M variant, the middle layers have a more pronounced influence. Overall, this provides a generalized view of the model’s behavior, serving as a sort of “fingerprint” for the model. Annex 9 provides an analysis for more model sizes and weights, as well as a visualization of the gradient distribution over all layers independently for each class in PascalVOC.
Figure 7: Gradient distribution over layers on PascalVOC for different pretrained ViT-L/14.
4.8 Qualitative Comparison to SOTA
Here, we present a qualitative analysis of the explainability maps generated by LeGrad in comparison to other SOTA methods. The results are depicted in Figure 6, which includes a diverse set of explainability approaches such as gradient-based methods (e.g., CheferCAM, GradCAM), attention-based methods (e.g., Raw Attention weights visualization, Rollout), and methods that integrate intermediate visual representations with text prompts (e.g., TextSpan). Our observations indicate that raw attention visualizations tend to highlight a few specific pixels with high intensity, often associated with the background rather than the object of interest. This pattern, consistent with findings in the literature [5, 11], suggests that certain tokens disproportionately capture attention weights. Consequently, methods that rely on raw attention weights to construct explainability maps, such as CheferCAM, exhibit similar artifacts. For instance, in the localization of ”Basket” (Figure 6, row 1), the basket is marginally accentuated amidst a predominance of noisy. In contrast, for LeGrad, the presence of uniform noisy activations across different prompts results in minimal gradients for these regions, effectively filtering them out from the final heatmaps. This characteristic enables LeGrad to produce more focused and relevant visual explanations.
@@ 5 Conclusion
We proposed LeGrad as an explainability method that highlights the decision-making process of ViTs across layers and for various feature aggregation strategies. We validated the method’s effectiveness in generating interpretable visual explanations that align with the model’s reasoning. Our approach offers a granular view of feature formation sensitivity, providing insights into the contribution of individual layers. The evaluation across object segmentation, perturbation, and open-vocabulary scenarios underscores LeGrad’s versatility and fidelity in different contexts. By facilitating a deeper understanding of ViTs, LeGrad paves the way for more transparent and interpretable models.
@@ 6 Overview
In this supplementary material, we first provide in section 7 detailes regarding the implementation, including pre-trained weights, baseline methodologies, and necessary adaptations for applying GradCAM to Vision Transformers (ViT). An expanded evaluation of LeGrad’s performance, specifically on Vision Transformers trained with the ImageNet dataset, is documented in section 8. Section 8.2 offers a visual representation of the perturbation benchmarks utilized to assess the efficacy of various explainability approaches, along with additional qualitative examples in section 8.3 to further illustrate the capabilities of our method. Section 9 further explores the gradient distribution across layers for different models and datasets. Furthermore, section 10 conducts a sanity check using the FunnyBirds Co-12 setup to evaluate the robustness of our explainability method. Lastly, section 11 includes a disclaimer on the use of personal and human subject data within our research.
@@ 7.1 Pretrained weights
The experiments conducted in our study leverage a suite of models with varying capacities, including ViT-B/16, ViT-L/14, ViT-H/14, and ViT-bigG/14. These models are initialized with pretrained weights from the OpenCLIP library respectively identified by: "laion2b_s34b_b88k", "laion2b_s32b_b82k", "laion2b_s32b_b79k", and "laion2b_s39b_b160k". For the SigLIP method, we utilize the ViT-B/16 model equipped with the "webl" weights. For the ”gradient distribution over layers” graphs, Figure 7, we also used the pretrained weights from OpenAI [29] and MetaCLIP [38].
@@ 7.2 Detailed Description of Baselines
In this section, we provide an overview of the baseline methods against which our proposed approach is compared.
@@
with $d$ representing the token dimension, * denoting element-wise multiplication, and the superscript $+$ the ReLU operation. We empirically determined that applying GradCAM to layer 8 of ViT-B/16 yields optimal results.
mIoU(V7) Neg $\uparrow$(INet) Pos $\downarrow$(INet)
GradCAM w/ [CLS] 8.72 45.26 22.86
GradCAM w/o [CLS] 8.18 41.29 24.20
Multi-layer w/ [CLS] 9.51 45.31 22.54
LeGrad 48.38 52.27 13.97
Table 7: Evaluation of GradCAM with and without [CLS] token as well as with layer aggregation as proposed in equation 8 on ViT-B/16 for open-vocabulary detection(V7) and perturbation (ImageNet).
@@
7.3 Details on adapting GradCAM to ViT
As GradCAM was designed for CNNs without [CLS] tokens, we tried both alternatives, i.e. including/excluding [CLS] token, Tab. 7 showcases a comparison of including/excluding the [CLS] token in the gradient computation on ViT-B. We observe that including the [CLS] token result in a marginal improvement. Overall, we would consider both options valid. GradCAM was originally designed for CNNs, therefore needs some adaptation to work for ViT. In an effort to consolidate the baselines used in this paper, we tried different configurations. One of which is whether or not include the [CLS] token in the gradient computation or not. We note that both alternatives are aligned with the original design of the GradCAM and that this choice is a matter of implementation.
We found that including the [CLS] token was producing better numbers, we therefore used that choice. Indeed, Table 7 shows the results on all the benchmark for GradCAM w/ and w/o the [CLS] token included in the gradient computation and shows that not including it translate in a slight decrease. Moreover, including the [CLS] token in the gradient computation is the option that makes more sense, as it is the [CLS] that is used to compute the similarity with the text query. We also tried to use the layer aggregation used in LeGrad for GradCAM and provide an evaluation in Table 7. We apply the same layer aggregation as in LeGrad (see 8) showing a slight improvement on the V7 benchmark. We attribute the only modest improvement to the fact that layers in ViT have different activation value ranges. LeGrad uses only gradients to produce layer-wise explainability maps, thereby avoiding this issue.
@@ 7.4 Adaptation of Baseline Methods to Attentional Pooler
CheferCAM: Following the original paper [9] that introduces CheferCAM, we considered the attentional pooler as an instance of a ”decoder transformer” and applied the relevancy update rules described in equation (10) of that paper [9], (following the example of DETR). Since the attentional pooler has no skip connection we adapted the relevancy update rule not to consider the skip connection.
@@
Table 8: SOTA comparison on ViT-B/16 on perturbation-based tasks on ImageNet-val for a ViT trained on ImageNet.
L/14 H/14
[…]
47.81 20.80 57.57 21.73
49.32 19.95 59.55 19.50
52.01 16.80 60.28 18.26
54.48 15.23 61.72 18.26
Table 9: Ablation study: ”ReLU” corresponds to whether or not negative gradients are set to 0. ”All layers” corresponds to whether or not the intermediate tokens are used to compute the gradient for every layer or if only the features from the last layer are used. Numbers are the AUC score for the perturbation base benchmark using the target class to compute the explainability map.
7.5 Mitigation of sensitivity to irrelevant regions:
$E^{s}_{E^{empty}>th}=0,$ (13)
where $th=0.8$. This method effectively addresses the issue without resorting to external data on the image content.
8 Additional results
8.1 Image Classification
In this section, we extend our evaluation of the proposed LeGrad method to Vision Transformers (ViTs) that have been trained on the ImageNet dataset for the task of image classification. The results of this evaluation are presented in Table 8, also providing a comparison with other state-of-the-art explainability methods.
Another observation is that even elementary explainability approaches, such as examining the raw attention maps from the final attention layer of the ViT, demonstrate competitive results. In fact, these basic methods surpass more complex ones like GradCAM (achieving an AUC of 53.1 versus 43.0 for negative perturbations).